OpenStudy (anonymous):

Solve: 9x^2 - 16x + 60 + 0

OpenStudy (luigi0210):

Quadratic formula?

OpenStudy (luigi0210):

Seems difficult to factor mentally.

OpenStudy (anonymous):

Yet agian

OpenStudy (ranga):

In your previous post you mentioned you were asked to find the discriminant first. Is this the same question?

OpenStudy (anonymous):

Kinda But i really just need the answer or to have it explaned to where i get the answer because its been 8 hours i have been working on this test.

OpenStudy (anonymous):

I would rather it be explained.

OpenStudy (ranga):

\[\Large x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a } = \frac{ -(-16) \pm \sqrt{(-16) ^{2} - 4(9)(60)} }{ 2(9) } \]

OpenStudy (anonymous):

whoa

OpenStudy (ranga):

\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a } = \frac{ -(-16) \pm \sqrt{(-16) ^{2} - 4(9)(60)} }{ 2(9) }\]

OpenStudy (anonymous):

Is that the quadratic formula?

OpenStudy (ranga):

Yes, that is the quadratic formula. Use your calculator and simplify the right hand side.

OpenStudy (anonymous):

ok one second than

OpenStudy (whpalmer4):

@Luigi0210 it seems difficult to factor mentally because it is irreducible :-)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got some crazy numbers

OpenStudy (anonymous):

16 √ 256 - 2,160 / 18 right?

OpenStudy (anonymous):

Mr @ranga did i do it correctly?

OpenStudy (ranga):

\[\frac{ 16 \pm \sqrt{256-2160} }{ 18 } = \frac{ 16 \pm \sqrt{-1904} }{ 18 }\]

OpenStudy (ranga):

Simplify further. There is a negative inside the square root. So this is a complex number. square root of -1 is "i". Also write 1904 as 119 * 16. square root of 16 is 4 which can be pulled out of the square root along with "i".

OpenStudy (anonymous):

-43.63 correct

OpenStudy (ranga):

\[\frac{ 16 \pm \sqrt{-1904} }{ 18 } = \frac{ 16 \pm \sqrt{-1 * 16 * 119} }{ 18 } = \frac{ 16 \pm 4i\sqrt{119} }{ 18 }\]

OpenStudy (anonymous):

could this just be 16-43.63/18 and 16 + 43.63/18 and simplyfy

OpenStudy (ranga):

\[= \frac{ 8 \pm 2i\sqrt{119} }{ 9 }\]

OpenStudy (ranga):

Unless they want the answer in decimal, leave it as a fraction. \[x = \frac{ 8 }{9} + \frac{2i\sqrt{119} }{ 9 }~~or~~x = \frac{ 8 }{9} - \frac{2i\sqrt{119} }{ 9 }\]

OpenStudy (anonymous):

Well is algebra 1? would it be a decimal?

OpenStudy (ranga):

Usually in math we leave it as fractions. In Science and Engineering applications we may resort to decimals.

OpenStudy (ranga):

But follow what your instructor says or see some class room examples.

OpenStudy (anonymous):

oh. ok. It seems when i have done this before i could leave it a fraction. P.S forgot everything.

OpenStudy (whpalmer4):

Yes, the "symbolic" form is much better than writing down decimals — often you can't even write the exact answer as a decimal! This would be one of those cases.

OpenStudy (whpalmer4):

Usually easy to crank out a number if you need it, but much harder to go from a number to an expression.

OpenStudy (ranga):

So that is the answer above. If you want you cam factor out 2/9 and write the answer as: \[x = \frac{ 2 }{9} (4 + i\sqrt{119})~~or~~x = \frac{ 2 }{9} (4 - i\sqrt{119})\]

OpenStudy (anonymous):

thats for the Fraction answer.?

OpenStudy (anonymous):

Ok thank you i have like 5 of these to do and this will help.

OpenStudy (ranga):

Alright. And good luck. Apply the quadratic formula and be careful when simplifying.

OpenStudy (ranga):

And yes, I meant "symbolic" and fractional form as whpalmer4 mentioned. The symbol here being the square root symbol. The fractional part is 2/9. 119 is not a perfect square and so the square root of 119 will be an irrational number. Rather than using the calculator to find square root of 119 we leave it in the symbolic form sqrt(119). Also we leave the fraction 2/9 as it is instead of concerting it into an approximate decimal.